# Sanjay Pradeepa, Chen Wangb, Matthew M. Dahmc, Jeff D. Eldredgea, and Candace S. J. Tsaid

### v1

@@ 4 Conclusions
Several promising extensions could further enhance the framework’s capabilities and applicability. First, incorporating classical morphological features such as aspect ratio, circularity, solidity, texture based features and Hu moments, alongside DINOv2 embeddings could improve discrimination of particles with subtle shape differences21, particularly for distinguishing loosely bundled clusters from compact matrix structures. Such hybrid feature representations would combine the semantic understanding of foundation models with explicit geometric constraints validated across decades of materials characterization. Second, the current mask-guided approach assumes that features within a masked region belong to a single morphological class, which becomes problematic when different particle types overlap. In such cases, the extracted features contain mixed signals from multiple morphologies, potentially confusing the classifier. To resolve this, future work could incorporate instance segmentation to detect and individually crop overlapping particles 14, 18. By upscaling and treating each crop as a standalone input, DINOv2 could classify the distinct morphology of each component independently, enabling robust analysis even in crowded fields. Third, integrating multi-modal data such as combining morphological classification from TEM with elemental composition from energy-dispersive X-ray spectroscopy (EDS) or structural information from electron diffraction, could provide more comprehensive particle characterization for complex engineered nanomaterials1.
The most impactful future direction involves retraining foundation models on large-scale, domain-specific electron microscopy datasets spanning diverse nanomaterial types, imaging conditions, and morphological variations. While transfer learning from natural image pre-training proves surprisingly effective, models trained explicitly on a large number of annotated EM images would likely capture nanoscale structural motifs, microscopy-specific artifacts, and fine morphological distinctions that current general-purpose vision models cannot reliably encode. Such domain-specialized foundation models—analogous to how medical imaging has developed radiology-specific vision transformers32—would represent a critical step toward fully automated quantification and classification of complex-shaped nanomaterials including CNTs, CNFs, metal oxide nanoparticles, graphene, and hybrid nanocomposites. Combined with advances in automated microscopy hardware and real-time image acquisition, these AI-enabled characterization pipelines could fundamentally transform nanomaterial quantification and classification from an art requiring expert manual oversight into a data-driven science.
@@
S.P. developed the methodology, performed all computational experiments, and wrote the manuscript. C.W. performed TEM analysis and image annotation. M.M.D. collected field samples and performed TEM analysis. J.D.E. provided supervision and manuscript review. C.S.J.T. provided domain expertise and manuscript review.
@@ Acknowledgements
  • M. E. Birch, C. Wang, J. E. Fernback, H. A. Feng, Q. T. Birch, and A. K. Dozier (2017) Analysis of carbon nanotubes and nanofibers on mixed cellulose ester filters by transmission electron microscopy. In NIOSH Manual of Analytical Methods (NMAM), Note: Chapter CN Cited by: §1.
@@
  • National Institute for Occupational Safety and Health (NIOSH) (2013) Current intelligence bulletin 65: occupational exposure to carbon nanotubes and nanofibers. Technical report Technical Report DHHS (NIOSH) Publication No. 2013-145, U.S. Department of Health and Human Services, Centers for Disease Control and Prevention, National Institute for Occupational Safety and Health, Cincinnati, OH, USA. External Links: Link Cited by: §1, §2, §2.
@@
  • A. Shah, J. A. Schiller, I. Ramos, J. Serrano, D. K. Adams, S. Tawfick, and E. Ertekin (2023) Automated image segmentation of scanning electron microscopy images of graphene using u-net neural network. Materials Today CommunicationsComputational Materials Sciencenpj Computational MaterialsarXiv preprint arXiv:1409.1556Nature MethodsarXiv preprint arXiv:2304.07193Journal of Nanoparticle ResearchRSC Adv.Frontiers in OncologyBeilstein Journal of NanotechnologyBioinformaticsnpj Computational MaterialsAmerican Industrial Hygiene Association JournalJournal of Toxicology and Environmental Health, Part B: Critical ReviewsJournal of Aerosol ScienceAnalytica Chimica ActaEnvironmental Science and TechnologyAnnals of Occupational HygieneConstruction and Building MaterialsComputational Materials ScienceActa MaterialiaIntegrating Materials and Manufacturing InnovationScienceJOMComputational Materials ScienceProgress in Materials ScienceApplied Surface ScienceMRS CommunicationsIEEE Transactions on Knowledge and Data EngineeringInternational Journal of Computer VisionCommunications of the ACMnpj 2D Materials and ApplicationsComputational Materials ScienceACS Central ScienceMRS BulletinAnalytical MethodsNatureAerosol Science and TechnologyJournal of Artificial Intelligence ResearchJournal of Big DataPattern Recognition LettersJournal of Computer and System SciencesComputational Statistics and Data AnalysisJournal of Machine Learning ResearchJournal of Machine Learning ResearcharXiv preprint arXiv:1301.3342Philosophical Transactions of the Royal Society A: Mathematical, Physical and Engineering SciencesJournal of Machine Learning ResearchPeerJ 35, pp. 106127. External Links: ISSN 2352-4928, Document, Link Cited by: §1.

### v2

@@ 4 Conclusions
Several promising extensions could further enhance the framework’s capabilities and applicability. First, incorporating classical morphological features such as aspect ratio, circularity, solidity, texture based features and Hu moments, alongside DINOv2 embeddings could improve discrimination of particles with subtle shape differences24, particularly for distinguishing loosely bundled clusters from compact matrix structures. Such hybrid feature representations would combine the semantic understanding of foundation models with explicit geometric constraints validated across decades of materials characterization. Second, the current mask-guided approach assumes that features within a masked region belong to a single morphological class, which becomes problematic when different particle types overlap. In such cases, the extracted features contain mixed signals from multiple morphologies, potentially confusing the classifier. To resolve this, future work could incorporate instance segmentation to detect and individually crop overlapping particles 17, 21. By upscaling and treating each crop as a standalone input, DINOv2 could classify the distinct morphology of each component independently, enabling robust analysis even in crowded fields. Third, integrating multi-modal data such as combining morphological classification from TEM with elemental composition from energy-dispersive X-ray spectroscopy (EDS) or structural information from electron diffraction, could provide more comprehensive particle characterization for complex engineered nanomaterials1.
The most impactful future direction involves retraining foundation models on large-scale, domain-specific electron microscopy datasets spanning diverse nanomaterial types, imaging conditions, and morphological variations. While transfer learning from natural image pre-training proves surprisingly effective, models trained explicitly on a large number of annotated EM images would likely capture nanoscale structural motifs, microscopy-specific artifacts, and fine morphological distinctions that current general-purpose vision models cannot reliably encode. Such domain-specialized foundation models—analogous to how medical imaging has developed radiology-specific vision transformers36—would represent a critical step toward fully automated quantification and classification of complex-shaped nanomaterials including CNTs, CNFs, metal oxide nanoparticles, graphene, and hybrid nanocomposites. Combined with advances in automated microscopy hardware and real-time image acquisition, these AI-enabled characterization pipelines could fundamentally transform nanomaterial quantification and classification from an art requiring expert manual oversight into a data-driven science.
@@
S.P. developed the methodology, performed all computational experiments, and wrote the manuscript. C.W. performed TEM analysis and image annotation. M.M.D. collected field samples. J.D.E. provided supervision and manuscript review. C.S.J.T. provided domain expertise and manuscript review.
@@ Acknowledgements
  • P. A. Baron (1994) Asbestos and Other Fibers by PCM: Method 7400. Issue 2 Technical Report Method 7400, National Institute for Occupational Safety and Health (NIOSH). Note: In NIOSH Manual of Analytical Methods (NMAM), 4th Edition External Links: Link Cited by: §2.
@@
  • M. E. Birch, C. Wang, J. E. Fernback, H. A. Feng, Q. T. Birch, and A. K. Dozier (2017) Analysis of carbon nanotubes and nanofibers on mixed cellulose ester filters by transmission electron microscopy. In NIOSH Manual of Analytical Methods (NMAM), Note: Chapter CN Cited by: §1, §1, §2, §2.
@@
  • M. M. Dahm, M. K. Schubauer-Berigan, D. E. Evans, M. E. Birch, S. Bertke, J. D. Beard, A. Erdely, J. E. Fernback, R. R. Mercer, and S. A. Grinshpun (2018) Exposure assessments for a cross-sectional epidemiologic study of US carbon nanotube and nanofiber workers. 221 (3), pp. 429–440 (en). Cited by: §2.
  • M. M. Dahm, D. E. Evans, S. Bertke, M. E. Birch, B. Ku, and K. H. Dunn (2015) Carbon nanotube and nanofiber exposure assessments: an analysis of 14 site visits. 59 (6), pp. 705–723. External Links: Document Cited by: §2.
@@
  • National Institute for Occupational Safety and Health (NIOSH) (2013) Current intelligence bulletin 65: occupational exposure to carbon nanotubes and nanofibers. Technical report Technical Report DHHS (NIOSH) Publication No. 2013-145, U.S. Department of Health and Human Services, Centers for Disease Control and Prevention, National Institute for Occupational Safety and Health, Cincinnati, OH, USA. External Links: Link Cited by: §1.
@@
  • P. A. Schulte, V. Leso, M. Niang, and I. Iavicoli (2019) Current state of knowledge on the health effects of engineered nanomaterials in workers: a systematic review of human studies and epidemiological investigations. Scand. J. Work Environ. Health 45 (3), pp. 217–238 (en). Cited by: §1.
  • A. Shah, J. A. Schiller, I. Ramos, J. Serrano, D. K. Adams, S. Tawfick, and E. Ertekin (2023) Automated image segmentation of scanning electron microscopy images of graphene using u-net neural network. Materials Today CommunicationsComputational Materials Sciencenpj Computational MaterialsarXiv preprint arXiv:1409.1556Nature MethodsarXiv preprint arXiv:2304.07193Journal of Nanoparticle ResearchRSC Adv.Frontiers in OncologyBeilstein Journal of NanotechnologyBioinformaticsnpj Computational MaterialsAmerican Industrial Hygiene Association JournalJournal of Toxicology and Environmental Health, Part B: Critical ReviewsJournal of Aerosol ScienceAnalytica Chimica ActaEnvironmental Science and TechnologyAnnals of Occupational HygieneConstruction and Building MaterialsComputational Materials ScienceActa MaterialiaIntegrating Materials and Manufacturing InnovationScienceJOMComputational Materials ScienceProgress in Materials ScienceApplied Surface ScienceMRS CommunicationsIEEE Transactions on Knowledge and Data EngineeringInternational Journal of Computer VisionCommunications of the ACMnpj 2D Materials and ApplicationsComputational Materials ScienceACS Central ScienceMRS BulletinAnalytical MethodsNatureAerosol Science and TechnologyJournal of Artificial Intelligence ResearchJournal of Big DataPattern Recognition LettersJournal of Computer and System SciencesComputational Statistics and Data AnalysisJournal of Machine Learning ResearchJournal of Machine Learning ResearcharXiv preprint arXiv:1301.3342Philosophical Transactions of the Royal Society A: Mathematical, Physical and Engineering SciencesJournal of Machine Learning ResearchPeerJInt. J. Hyg. Environ. Health 35, pp. 106127. External Links: ISSN 2352-4928, Document, Link Cited by: §1.